### v1

@@ Abstract
Figure 3: (color online): (a) Band structures of $H(k)$. After cyclic TPT, stable population is established near gap closing (near $\Gamma$, orange region). (b) The $k$ is interpreted as the coordinate of a continuous path in $({\Theta},{\Phi})$ space ($S^{2}$ surface shown by the inset). Every time longitude ${\Phi}=0$ (dashed) is passed, pumping probability $p_{G}$ gets to maximum (1st, 2nd, 3rd…). The color scales of dots stand for $k$’s value in unit of $\pi$. (c) Analytic solutions of $p_{G}$ (for spin 1/2). (d) Numerical results of $p_{G}$ in B.Z. ($p_{G}(k)=p_{G}(-k))$ subject to different phonon period ${\tau}_{ph}$, and its correspondence to $({\Theta},{\Phi})$ (lower panels). Maximum of $p_{G}$ occurs at dashed line (lower panels) crossing $\Phi$ (orange dots). Since $\Phi$ is periodic, the “jump” from -1 to 1 is actually continuous. (e) Oscillation of pumping tuned by bias voltages.
Conclusion. Our goals are twofold. First, elucidate the origin of geometry, endowing its entrance with a generic source, unconstrained by specific scenarios. In view of existing proposals involving different levels of quantum, statistics, etc., 2 ; 3 ; 4 ; 7 ; 8 ; 9 we are not expecting the framework is universal, but it might furnish a more unified one. Second, show the existence and influence of quantum geometry with observables in spin and band. Intuitively speaking, topology is demonstrated by “steps”, and we find geometry could be demonstrated by some “wavy lines”; moreover, these wavy lines can be tuned (for spin, by the loop vertex angle $\Theta$, and for band by band inversions, i.e., the presence or absence of TPT 26 ; 29 ), such that their origin could be tested and confirmed. The technique $Z$-map is useful for assigning a geometric interpretation, e.g., oscillation is circling around $S^{2}$ and dephasing is approaching to the poles. It is also useful in finding analytic solutions Eqns. 6,  7.

### v2

@@ Abstract
Figure 3: (color online): (a) $k$ serves as the coordinate of a path in $({\Theta},{\Phi})$ space ($S^{2}$ surface shown by the inset). Every time longitude ${\Phi}=0$ (dashed) is passed, pumping probability $p_{G}$ gets to maximum (1st, 2nd, 3rd…). The color scales of dots stand for $k$’s value in unit of $\pi$. (b) Analytic solutions of $p_{G}$ (for spin 1/2). (c) Numerical results of $p_{G}$ in BZ. ($p_{G}(k)=p_{G}(-k))$ subject to different phonon period ${\tau}_{ph}$, and its correspondence to $({\Theta},{\Phi})$ (lower panels). Maximum of $p_{G}$ occurs at dashed line (lower panels) crossing $\Phi$ (orange dots). Since $\Phi$ is periodic, the “jump” from -1 to 1 is actually continuous. (d) Oscillation of pumping tuned by bias voltages.
Outlook. $Z$-map proves useful in revealing anomalies and a deep-seated contradiction: the monotonic dependence of pumping on gaps or frequencies, while intuitively satisfying, secretly violates the principles of unitarity and continuity. $Z$-map is independent of energetic parameters and space dimensions, and its strength is to uncover generic behaviors across models. Therefore, a promising direction is by judicious construction of different $Z$-maps (Fig. 1) to discover more such behaviors, analogous to on-going efforts in expressing different quantities in terms of quantum metrics a4 ; a5 ; a6 ; a7 . However, the focus shifts to general geometry rather than a specific geometric quantity. Another inspiration of $Z$-map is that geometry’s influence can go beyond analytic ways, as exemplified by the connection between dephasing and “poles”.
Conclusion. First, geometry enters quantum via unitarity, unconstrained by specific approximations or scenarios, leading to generic phenomena and rigorous concepts. Second, observables are uncovered for visualizing the influence of quantum geometry in spin/band scenarios. Intuitively speaking, topology is demonstrated by “steps”, and here geometry can be demonstrated with “wavy lines”; moreover, these wavy lines are abnormal to conventional belief and can be tuned (for spin, by the loop vertex angle $\Theta$, and for band by band inversion 29 ; 32 ), such that their identities could be tested.
I Appendix
I.1 A. Generality of $Z$-map & Its limitation
(1) Z-map Unifies different models in different spaces. The unifying comes in from several tiers. First, different models, say $({\Theta},{\omega}_{1}),({\Theta},{\omega}_{2}),...$ in the original $X_{e}$, described by the standard “pie-slice” loops (Fig. 2a) could be mapped to a common point in $X_{g}$. That is, a point in geometric space $X_{g}$, say $({\Theta},{\Phi})$, is not just standing for a single model, but an equivalence class of models that lead to a common evolution operator. In other words, the unifying is related to the fact that $Z$-map is non-injective (multiple-to-one). The benefit is that, to examine “all” models, we not necessarily exhaustively write down all Hamiltonians, but only select one representative in each equivalence class, such as using the “pie-slice” model. So, instead of covering all possible Hamiltonians, we try to cover all evolution outcomes of Hamiltonians.
Second, following this idea, the unifying could go beyond the standard loop to arbitrary models, as long as they are characterized by a set of parameters. For example, a “bumpy” loop (Fig. 2a) expands $X_{e}$ from $(B,{\omega})$ to a higher dimension ${\{}B_{j},{\omega}_{j}{\}}$ (Fig. 2 and SI). Notably, increasing $X_{e}$ will not change $X_{g}$. The idea is that a model ${\{}B_{j},{\omega}_{j}{\}}$ in the expanded $X_{e}$ will still find a place on the same $X_{g}$. Formally, this is expressed with a map “distortion” $({\Theta},{\Phi}){\mapsto}({\Theta}^{\prime},{\Phi}^{\prime})$ (Fig. 2a). That means arbitrary models characterized by a specific $X_{e}$ could be unified to a common $X_{g}$.
Third, different scenarios, such as spin, band, which are formulated by different $X_{e}$ spaces, could still be mapped to a common $X_{g}$. Moreover, the dimension of $X_{e}$ is independent of the dimension of Hilbert space. In the paper, we show with spin 1/2 and spin 1, which reside in 2D and 3D Hilbert spaces, while they share a common $X_{e}=(B,{\omega})$. By this example, we see arguments, such as dephasing at ${\Theta}{\mapsto}{\pi}$, can be made across different Hilbert spaces. Otherwise, we would have faced an infinite task of studying different dimensions one by one.
In short, different Hilbert spaces are unified to a common $X_{e}$; different $X_{e}$ could be unified under a common $X_{g}$; then we work on $X_{g}$ — This is the “mechanism” for making generic arguments across scenarios, models, and Hilbert spaces.
(2) General applicability and phenomena. Berry phase involves (i) geometric terms (connection, curvature, etc.) (ii) in generic setups. Notably, most protocols for quantum geometry only inherit (i) by exploring further geometric term, such as metric; but lose (ii), because to make metric enter, it relies on specific approximations or scenarios. By contrast, Berry phase is defined in generic parameter space, independent of whether the system is atom or crystal, whether the dispersion is flat or not.
$Z$-map tries to maintains both (i) and (ii). Firstly, on $X_{g}$, various geometric quantities (including metric) could be defined; secondly, the existence of $X_{g}$ merely relies on unitary, thus (ii) is respected too.
Generality is also seen from the general observables uncovered. It is often said quantized ${\sigma}_{xy}$ is general, because it is due to “continuity”, which forbids the wavefunction jumps between topologies. This generality makes the effect robust, insensitive to system’s details.
Here, $Z$-map predicts robust phenomena, such as geometric oscillation/dephasing, which also arise from “continuity”. Such phenomena might take place in different scenarios and parameter regimes. For example, we show the robustness with different phonon periods ${\tau}_{ph}$ (independent of phonon modes). Besides, the robustness is demonstrated by parameter-independent phenomena, such as geometric oscillation of half-integer $s$ along ${\omega}$ (or $f$) being twice faster than integer $s$ (Fig. 2e, f),
(3) Relationship with metric-based geometry. $Z$-map is a comprehensive framework for various geometric quantities to be built in, such as metric (or particularly Fubini-Study metric), measure, curvature. The entrance is just $X_{g}$ space. Moreover, $X_{g}$ (solely) arises from unitary evolution, thus naturally underscores the Fubini-Study metric as a unitary invariant. Otherwise, if the formulation merely involves gauge invariant, Fubini-Study metric does not prove irreplaceable, thus not “physical”.
Besides, $Z$-map extends the ways of geometry giving the influence. Most commonly, geometry enters locally, in forms of a local field, such as metric 7 ; 8 ; 9 ; 18 ; 34 ; 35 ; 36 ; 37 , curvature ; or sometimes enters in a global way by an integration of metric 19 ; 24 . The geometric oscillation/dephasing from $Z$-map are both due to global geometry, but not due to an integration of an explicit local field. Besides, it raises the possibility of non-analytic ways of geometry giving its influence, such as the poles when one space tries to cover a different one.
Limitation of Z-map. $Z$-map generalizes geometry from metric to measure, and possibly to other aspects. But $Z$-map will not unify every existing scheme as special cases. Because their “inputs” could be different. $Z$-map is stricter in generality in introducing geometric concepts. In short, they will overlap, but totally inclusive.
Another limitation is that $Z$-map is only to reveal generic behaviors, and detailed information will be ignored. For example, a generic oscillation can be proved by geometry, while the exact locations of peaks or valleys are scenario specific. This is in analog with quantized ${\sigma}_{xy}$, which is due to edge bands protected by topology, thus material-independent; but the detailed dispersion and locations of those edge bands are material-dependent.
I.2 B. Viability of experiments
First, the observable is an oscillation pattern, more detectable than fine features. Second, unlike some fragile phenomena (e.g., spin liquid) living in narrow regimes, the oscillation is robust to model details, easier for material selection and preparation. Third, it is scenario independent, that means it allows crosscheck in different systems from local spins to various narrow-gapped topological insulators, along different axes, such as $\omega$, ${\epsilon}_{0}$(Fig. 2,3). Fourth, the frequency of $\textbf{B}(t)$ is MHz, and the driving of bands is in THz — both are technically achievable.
In particular, for the band model, we already made preliminary measurement in ZrTe5 32 ; 33 . The TPT could be realized by $A_{1g}$, $B_{1u}$ phonon modes, which can be excited by THz or optical pulses. The pumping physically is a meta-stable charge configuration after a relatively “long” time, thus, it is insensitive to the vibration process, as long as the phonon will close up the gap for a number of times. The charge pumping can be measured by transition rate change with ${\sim}10^{-5}$-$10^{-6}$ and sub-ps time resolution. Next, we need to extend the measurement under a biased voltage or applied strains.
Parameters used for the experimental proposal. In Fig. 2e, f. The frequency $f$ is $1/{\tau}$, where ${\tau}$ is the period of $\textbf{B}(t)$ completing a whole loop. Magnetic resonance experiment is typically performed in a MHz range 38 . Thus, parameters will be chosen to approach that range. By a rough estimate, $\bar{B}=0.01$ T is suitable and technically achievable. Assume ${\Phi}_{0}=0$ at $f_{0}=1/{\tau}_{0}=10$ MHz for a demonstration. (Otherwise, if ${\Phi}=0$ happens at a different frequency, the oscillation simply has an overall shift to the true $f_{0}$.) Then, ${\Phi}={\Phi}_{0}+{\Delta}{\Phi}=\bar{E}{\cdot}({\tau}-{\tau}_{0})/{\hbar}$ and $\bar{E}={\mu}_{B}\bar{B}$, where $\bar{E}$ and $\bar{B}$ are the average energy and magnetic field. Some calculations give ${\Phi}=c(1/f-{\tau}_{0})$, where $c$ is a coefficent ${\sim}2.8{\pi}{\cdot}10^{2}$ MHz. Combining $\Phi$ with Eqs. 6,7, we may plot the oscillations under different $\Theta$.
As a necessary check, we should remember $f$ cannot be too fast (or equivalently $\bar{B}$ cannot be too small), because the pumping is derived based on $hf{\ll}{\bar{\Delta}}$. In this case, $\bar{B}=0.01$ T leads to $\bar{\Delta}{\sim}10^{-27}$ J and $hf$ is ${\sim}10^{-25}$ J. Thus, the precondition is satisfied. In other words, it is well separated from a Rabi pumping 31
In Fig. 3, we have adopted $c_{H}=0.5$ eV, which gives a Fermi velocity under this model ${\sim}10^{5}$ m/s, consistent with ZrTe5. More parameters, such as the sampling density in BZ, could be found in SI.